OpenStudy (anonymous):

@ganeshie8 can you help me? i'll post the question below, it's a quadratic formula question...

OpenStudy (anonymous):

\[3x ^{2}+7x-24+13x\]

OpenStudy (anonymous):

Use the quadratic formula to solve and use factoring to check your answer (I already know the answer is -2,4 but i keep getting 1,3

OpenStudy (anonymous):

that should be = 13x not +13x

Directrix (directrix):

@ilikequadraticsokay Please close this thread given that it is the same question which you posted. Thanks. http://openstudy.com/study#/updates/529e87dce4b0ad25f2dee29a

OpenStudy (ranga):

3x^2 + 7x - 24 = 13x 3x^2 + 7x - 13x - 24 = 0 3x^2 - 6x - 24 = 0 divide by 3 x^2 - 2x - 8 = 0 discriminant = b^2 - 4ac = (-2)^2 - (4)(1)(-8) = 4 + 32 = 36 roots = { -(-2) +/- sqrt(36) } / 2 = (2 +/- 6) / 2 = 1 +/- 3 = -2 or 4

OpenStudy (anonymous):

i dont know what the discriminant is

OpenStudy (ranga):

x^2 - 2x - 8 = 0 factoring x^2 - 4x + 2x - 8 = 0 x(x - 4) + 2(x - 4) = 0 (x - 4)(x + 2) = 0 x = -2 or +4

OpenStudy (ranga):

just the term that goes inside the square root in the quadratic formula: b^2 - 4ac is called the discriminant.

OpenStudy (anonymous):

Oh okay, one moment.

OpenStudy (anonymous):

How did you go from 1 +/- 3 to -2,4?

OpenStudy (ranga):

1 + 3 = 4 1 - 3 = -2

OpenStudy (anonymous):

That makes sense.

OpenStudy (ranga):

alright.

OpenStudy (anonymous):

So I was just forgetting to actually solve the +/- part?

OpenStudy (ranga):

not sure. if you retrace your steps you may be able to spot the mistake.

OpenStudy (ranga):

x^2 - 2x - 8 = 0 \[\Large \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a } = \frac{ -(-2) \pm \sqrt{(-2) ^{2} - 4(1)(-8)} }{ 2(1) }\]

OpenStudy (ranga):

\[= \frac{ 2 \pm \sqrt{4 + 32} }{ 2 } = \frac{ 2 \pm \sqrt{36} }{ 2 } = \frac{ 2 \pm 6}{ 2 } = 1 \pm 3\]

OpenStudy (ranga):

1 + 3 = 4 1 - 3 = -2

OpenStudy (anonymous):

Yeah, I did all that. I guess I just thought it ended at 1 +/- 3

OpenStudy (ranga):

oh, ok. you will find two roots: one by adding and another by subtracting the second number.

OpenStudy (anonymous):

Okay, thank you so much!

OpenStudy (ranga):

you are welcome.